If a radius of a certain circle is tripled, with 8.2 centimeters then added,
the result is the circumference of the circle. Find the radius of the circle.

Answer by Theo(13342)

r = radius of the circle.

circumference of the circle = 2*pi*r

3*r + 8.2 = 2*pi*r

solve for r.

subtract 2*pi*r from both sides of the equation and subtract 8.2 from both sides of the equation to get:

3*r - 2*pi*r = -8.2

factor out the r to get:

r * (3-2*pi) = -8.2

divide both sides of the equation by (3-2*pi) to get:

r = -8.2 / (3-2*pi)

simplify to get:

r = 2.49757453

3*r + 8.2 = 156.9272359

2*pi*r = 156.9272359

Since they are the same, the answer is confirmed.

the answer is:

r = 2.49757453 centimeters.